OpenStudy (anonymous):

find all the points of the form (x,-x) which are 1 unit from the origin

OpenStudy (anonymous):

i know to use the distance formula do i set the d=1?

Directrix (directrix):

x^2 + ( -x)^2 = 1^2 x² + x² = 1 and then? @ninab731 r = 1 because the circle has center at the origin and the points you are looking for are 1 unit from the origin and of the form (x,-x)

Directrix (directrix):

What does x² + x² equal when you add them?

OpenStudy (anonymous):

2x^2

Directrix (directrix):

So, 2*x² = 1 What does just x² equal?

Directrix (directrix):

Divide both sides of this --> 2*x² = 1 by 2. What is 2*x² divided by 2 ?

OpenStudy (anonymous):

x^2=1/2?

Directrix (directrix):

Correct.

Directrix (directrix):

Then, x = ± √(1/2). Does that need to be rationalized? Does your teacher prefer that?

OpenStudy (anonymous):

the answer is in the book as \[(\frac{ \sqrt{2} }{ 2},- \frac{ \sqrt{2} }{ 2 }), (-\frac{ \sqrt{2} }{ 2 },\frac{ \sqrt{2} }{ 2 }\]

Directrix (directrix):

v